Minute of KAGRA commissioning meeting
- Date: 2018/6/26 11AM~12PM
- Attendance: Miyakawa, Yokozawa, Stefan, Nakano, Yamamoto
- eZuce: Shoda, Ushiba, Hirose, Eleonora, Izumi, Yamamoto(Toyama), Kokeyama, Miyoki
- Summary of PMC work by Stefan during his stay at KAGRA
- Two PZT steering mirrors were installed yesterday for input beam to IMC.
- Calibration to Hz will be done today.
- Thermal drive is sill missing.
- Needs fitness, noise budget or something.
- Probably we do not need to change HV PZT driver by O3, but needs estimation.
PMC servo modifications -> Stefan
FSS and PMC locked together -> Nakano, Stefan
PMC autolocking sped up -> Stefan, Nakano
Beam height brain storming -> next week?
- ETMX and ITMX should be located at half low temperature stage and room temperature stage. It means that ~5mm off from the nominal position. Type A can compensate ~5mm.
Necessary number of High power coil driver -> Aso
- Cable cleaning on BS rack will be done this weekend.
- Next meeting: 7/3.